# Visual Acoustic Fields

### v1

@@ 4 Dataset Statistics
Table 1: We compare our dataset with existing hitting sound datasets in terms of scenario (object-level vs. scene-level), data source (synthetic, robotic, or human-collected), and whether the dataset contains 3D spatial information. Unlike prior datasets focusing on object-level interactions or synthetic environments, our dataset captures real-world, human-collected impact sounds at the scene level with full 3D spatial alignment.
@@ 5 Experiments
As shown in Table 4, we measure Acc(1) and Acc(3). Acc(1) represents the accuracy rate when the segment with the highest relevance score is the correct result, while Acc(3) indicates the accuracy rate when the correct location is covered by one of the top three segments with the highest relevance scores. Our method significantly outperforms the baselines in both top-1 and top-3 localization accuracy rates. This suggests that existing large-scale sound datasets, such as AudioSet [14], lack a sufficient number of clean hitting sound samples for models to effectively perform localization tasks, despite the prevalence of hitting sounds in everyday life. With our collected dataset and proposed pipeline, we can effectively localize sound in 3D.
@@
In Fig. 6, we present visualizations of our generated sound. The Mel spectrogram reveals that our generated sounds exhibit similar frequency content and temporal evolution, aligning with our survey results, which indicate that our generated sound is nearly indistinguishable from the ground truth sound.
@@ 5.3 Implementation Details
RGB and Feature Field. We follow the official 3DGS implementation [23] to render RGB images. For feature extraction, we use the Langsplat codebase [34] to obtain multilevel SAM features and reconstruct the feature field. During data collection, we capture a video for each scene and uniformly sample approximately 300 frames from it as the training set. We train both the RGB and feature fields on a single NVIDIA GeForce RTX 3090 GPU.
Sound Generation and Localization. For sound generation, our implementation builds upon Stable Audio Open [9], using its released checkpoint as the starting point for training. We modify its conditioning encoder to process visual features by incorporating MLPs. During training, we optimize the conditioning encoder using the AdamW optimizer with a base learning rate of $5e^{-5}$ and a batch size of 8 on a single NVIDIA A100 GPU. For inference, we apply classifier-free guidance with a scale of 6 and perform 250 sampling steps. To evaluate sound quality, we use the codebase from [29]. For sound localization, we train AudioClIP [15] on our dataset using an SGD optimizer with a learning rate of $5e^{-5}$ on an NVIDIA GeForce RTX 3090 GPU. We initialize the image encoder with pretrained CLIP [35] weights and train only the audio encoder.
Baseline. For the baseline, we use the open-sourced implementation of im2wav [41]. Since, by default, the model cannot handle very short sequences, such as our hitting sounds, we pad them to 10 seconds for training and inference, then crop them afterward for metric calculation. In im2wav, we use a batch size of 16 for the VQVAE and a batch size of 8 for the upsampling and low-level models. All baseline training is done on an NVIDIA GeForce RTX 3090 GPU.
@@ 6 Conclusion, Limitation, and Future Work
Limitations and Future Works. Despite the promising results, our approach has several limitations. First, our dataset, while diverse, remains limited to static scenes and a variety of materials. Expanding the dataset to include more environments (including dynamic scenes and more outdoor settings) and objects with richer acoustic properties would enable better generalization of our model. Another limitation of our Visual Acoustic Fields model is its agnosticism of the spatial location of the listener. Specifically, our model does not account for the perceived sound changes based on the listener’s position. To overcome this limitation, future research could incorporate physics-based sound propagation models that account for distance attenuation, occlusion, and echo.

### v2

@@ 4 Dataset Statistics
Table 1: We compare our dataset with existing hitting sound datasets in terms of scenario (object-level vs. scene-level), data source (synthesized by simulators, collected by robot, or human-collected), and whether the dataset contains 3D spatial information. Unlike prior datasets focusing on object-level interactions or synthetic environments, our dataset captures real-world, human-collected impact sounds at the scene level with full 3D spatial alignment.
@@ 5 Experiments
As shown in Table 4, we measure Acc(1) and Acc(3). Acc(1) represents the accuracy rate when the segment with the highest relevance score is the correct result, while Acc(3) indicates the accuracy rate when the correct location is covered by one of the top three segments with the highest relevance scores. Our method significantly outperforms the baselines in both top-1 and top-3 localization accuracy rates. This suggests that existing large-scale sound datasets, such as AudioSet [14], lack a sufficient number of clean and high quality hitting sound samples to train a contrastive learning model and effectively perform localization tasks, despite the prevalence of hitting sounds in everyday life. With our collected dataset and proposed pipeline, we can effectively localize sound in a 3D scene.
@@
In Fig. 6, we present visualizations of our generated sound, the baseline, and the groundtruth sound. The Mel spectrogram reveals that our generated sounds exhibit similar frequency content and temporal evolution compared with the groundtruth sound. This aligns with our survey results, which indicate that our generated sound is nearly indistinguishable from the ground truth sound.
@@ 5.3 Implementation Details
RGB and Feature Field. We follow the official 3DGS implementation [23] to render RGB images. For feature extraction, we use the Langsplat codebase [34] to obtain multilevel SAM features and reconstruct the feature fields. During data collection, we capture a video for each scene and uniformly sample approximately 300 frames from it as the training set for scene reconstruction. We train both the RGB fields and feature fields on a single NVIDIA GeForce RTX 3090 GPU.
Sound Generation and Localization. For sound generation, our implementation builds upon Stable Audio Open [9], using its released checkpoint as the starting point for training. We modify its conditioning encoder to process visual features by incorporating MLPs. During training, we optimize the conditioning encoder using the AdamW optimizer with a base learning rate of $5e^{-5}$ and a batch size of 8 on a single NVIDIA A100 GPU. For inference, we apply classifier-free guidance with a scale of 6 and perform 250 sampling steps. To evaluate sound quality, we use the codebase from [29]. For sound localization, we train AudioClIP [15] on our dataset using an SGD optimizer with a learning rate of $5e^{-5}$ on a single NVIDIA GeForce RTX 3090 GPU. We initialize the image encoder with pretrained CLIP [35] weights and freeze it during training. Only the weights of audio encoder part are updated.
Baseline. For the baseline, we use the open-sourced implementation of im2wav [41]. Since, by default, the model cannot handle very short sequences, such as our hitting sounds, we pad them to 10 seconds for training and inference, then crop them afterward for metric calculation. When training im2wav, we use a batch size of 16 for the VQVAE and a batch size of 8 for the upsampling and low-level models. All baseline training is done on a single NVIDIA GeForce RTX 3090 GPU.
@@ 6 Conclusion, Limitation, and Future Work
Limitations and Future Works. Despite the promising results, our approach has several limitations. First, our dataset, while diverse, remains limited to static scenes and a variety of materials. Expanding the dataset to include more environments, such as including dynamic scenes and more outdoor settings, and objects with richer acoustic properties would enable better generalization of our model. Another limitation of our Visual Acoustic Fields model is its agnosticism of the spatial location of the listener. Specifically, our model does not account for the perceived sound changes based on the listener’s position. To overcome this limitation, future research could incorporate physics-based sound propagation models that account for distance attenuation, occlusion, and echo.